abal is dropped from a height of 16m and rebund to 9m.if contact time is 0.2 sec. find acc.

Grade:11

1 Answers

Vikas TU
7 years ago
Writing the eqn. for ball we get,
16 = 0 + 0.5gt^2
t = root(32/g) s.= 1.788 s
v = 0 + 10*1.78 = 17.8 m/s (final velocity)
 
After rebounding,
0^2 = 17.8^2 – 2*(a-g)*9
solve for a.